OpenStudy (anonymous):

Find the domain of: y = 5/2x + 24 A: all real numbers except 12 B: all real numbers except 0.08 C: ll real numbers except –0.08 D: all real numbers except –12

OpenStudy (anonymous):

C is all real numbers vs ll ops xD

OpenStudy (anonymous):

all real except zero

OpenStudy (anonymous):

Can't pic that

OpenStudy (anonymous):

if u put zero in the equation will result in undefined so zero is not domain all others are domain

OpenStudy (anonymous):

i Don't see what you are tlakign about, there are no zeros

OpenStudy (anonymous):

ur choices are wrong, it does not fulfil the definition of domain

OpenStudy (anonymous):

they are my choices

OpenStudy (anonymous):

non of the given is correct, the correct one is all real numbers except zero

OpenStudy (anonymous):

let me post pic

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Pic just cuz i think i typed it fk'ed up -_-

OpenStudy (anonymous):

now ur question is correct i will ans in a mint

OpenStudy (anonymous):

all real numbers except -12

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thank you so what did i type rung?

OpenStudy (anonymous):

u typed y = 5/2x + 24 it must be y = 5/(2x + 24)